Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by NarendraSure » Sat Jun 27, 2015 8:50 pm
Molly needs to roll a 1 in a die and she has 3 tries. What is the probability that she will succeed?


can some body answer this with an explanation.

Regards
Naren
Join the discussion
Source: — Problem Solving |

by NarendraSure » Sat Jun 27, 2015 8:55 pm
Hi All,

Thanks for passing by, found the answer. If anyone is interested below is how.

Probability to roll a 1 at the first try: 1/6
Probability to roll a 1 at the second try: 5/6 * 1/6 = 5/36
Probability to roll a 1 at the third try: 5/6 * 5/6 * 1/6 = 25/216
Since she can succeed with one OR two OR three tries, we sum it: 91/216

P(2nd roll) is the probability that you DON'T roll a 1 on the 1st roll, because the only way to get to the 2nd roll is to avoid a 1 on the 1st roll. this is why we need to multiply the 1/6 by 5/6.



Regards
Naren
Join the discussion

by Brent@GMATPrepNow » Sun Jun 28, 2015 9:23 am
NarendraSure wrote:Molly needs to roll a 1 in a die and she has 3 tries. What is the probability that she will succeed?


can some body answer this with an explanation.

Regards
Naren
We want P(at least one 1 in three rolls).

When it comes to probability questions involving "at least," it's best to try using the complement.
That is, P(Event A happening) = 1 - P(Event A not happening)
So, here we get: P(getting at least one 1 in three rolls) = 1 - P(not getting at least one 1 in three rolls)
What does it mean to not get at least one 1? It means getting zero 1's.
So, we can write: P(getting at least one 1) = 1 - P(getting zero 1's in three rolls)

P(get zero 1's in three rolls)
P(get zero 1's in three rolls) = P(no 1 on 1st roll AND no 1 on 2nd roll AND no 1 on 3rd roll )
= P(no 1 on 1st roll) x P(no 1 on 2nd roll) x P(no 1 on 3rd roll)
= 5/6 x 5/6 x 5/6
= 125/216

So, P(getting at least one 1) = 1 - 125/216
= [spoiler]91/216[/spoiler]

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by NarendraSure » Mon Jun 29, 2015 4:22 am
Thanks for the reply Brent.

Regards
Naren
Join the discussion

by Matt@VeritasPrep » Mon Jun 29, 2015 3:26 pm
Just to illustrate why Brent's approach is so good, consider the positive approach, in which we list all the cases.

To get exactly one 1, we would need to roll 1, not-1, not-1 in some order. That gives us

1XX = (1/6) * (5/6) * (5/6) = 25/216

But this could happen in any of THREE orders: 1XX, X1X, XX1. So we have 3 * (25/216) = 75/216.

Now we consider exactly two 1s, which is basically the same idea:

11X = (1/6) * (1/6) * (5/6) = 5/216

This could happen in three orders as well (11X, 1X1, X11), so 3 * (5/216) = 15/216.

We also have exactly three ones, which is 111 = (1/6)*(1/6)*(1/6) = 1/216.

Adding these up, we have 75/216 + 15/216 + 1/216 = 91/216. Yuck!

So doing P(At least one) = 1 - P(No ones) is much more efficient.
Join the discussion

by NarendraSure » Wed Jul 01, 2015 5:50 am
Yeah agree matt this is so direct
Join the discussion

by talaangoshtari » Wed Jul 01, 2015 6:12 am
In probability, we have a distribution called Geometric Distribution.
P(X = x) = p.(q)^(x - 1)
p = probability of success
q = probability of failure

Here we have,
P(1《 X 《 3) = P(X = 1) + P(X = 2) + P(X = 3)
= p + pq + pq^2 = (1/6) + (1/6)(5/6) + (1/6)(5/6)^2 = 91/216
Join the discussion

by NarendraSure » Wed Jul 01, 2015 1:49 pm
Hi talaangoshtari,

how do we need to determine the usage of this neat little formula.

Regards
Naren
Join the discussion

by talaangoshtari » Wed Jul 01, 2015 9:55 pm
Hi NarendraSure,

Geometric distribution computes the probabability of the first success after k - 1 failures. In this problem there are 3 scenarios:

1. We roll a 1 in a die in the first try, so we have one success.
2. We roll a 1 in a die in the second try, so we have one failure and one success.
3. We roll a 1 in a die in the third try, so we have two failures and one success.
Join the discussion

by nikhilgmat31 » Fri Jul 03, 2015 12:33 am
Brent's way to solve is simple calculation. Great.

But it is difficult to strike. But it is very good to solve such questions

P(A) = 1- P(Not A)
Join the discussion